OpenStudy (anonymous):

Complete the Next Step: A partial solution to the equation 5(2x + 3) = 85 is given below. Choose the step that is missing from the solution.

OpenStudy (anonymous):

10x + 3 = 85 2x + 15 = 85 10x + 15 = 17 10x + 15 = 85

OpenStudy (anonymous):

Are you asking which of the above is the next step?

Parth (parthkohli):

I wonder why they don't try to make it easier by making you divide both sides by 5 instead of distributing the whole thing.

OpenStudy (anonymous):

10x + 15 = 85

Parth (parthkohli):

Distribute \(5(2x + 3)\)

OpenStudy (anonymous):

5*2x + 5*3 = 10x + 15 as @Ichigo23 and @ParthKohli have mentioned.